Judy's average on 4 tests is 80. What is the least that she can earn on her 5th test and still have a chance for an 85 average on her 7th test.
Answer by JulietG(1812)

The total number of points available is 7 tests @ 100 points, or 700.
She wants to have at least an 85 average after 7 tests.
7 * 85 = 595.
She already has 4 * 80 = 320.
Since you're looking for the least on the 5th, assume that she scores 100 on the 6th and 7th.
320 + 100 + 100 + x = 595
520 + x = 595
Subtract 520 from each side
x = 75
.
If Judy scores 100 on test 6 and test 7, she will need a 75 or higher on test 5 to come out with an 85 average.